I am trying to help my sister with the inverse variation. I am not sure how to do that type of problem. Can you please walk me through step by step?
thanks,
mw

Answer by stanbon(75887)

I am trying to help my sister with the inverse variation. I am not sure how to do that type of problem. Can you please walk me through step by step?
---------------------
If x and y vary inversely y = k/x expresses that fact.
-----------------------------------------
Assume that y = 12 when x = 2 and you want to find out
what y is when x = 10.
--------------------
1st: you have to find the value of "k".
You let y = 12 and x = 2 and solve for "k", as follows:
12 = k/2
k = 24
---------------------
Now you know that y = 24/x for this situation.
---
You still want to find out what y is when x = 10,
so you substitute and get
y = 24/10
You find that y = 2.4 when y = 10
=============================
Hope that helps.
Cheers,
Stan H.